the product of two consecutive odd integers is 77 more than twice the larger, Find the integers please. I cannot set up "product" consecutive integers?
Found 3 solutions by Alan3354, stanbon, laurien56:
Answer by Alan3354(69443)

the product of two consecutive odd integers is 77 more than twice the larger, Find the integers please. I cannot set up "product" consecutive integers?
---------------
The product is x*(x+2)
---
x*(x+2) = 2(x+2) + 77
x^2 + 2x = 2x + 4 + 77
x^2 - 81 = 0
x = ± 9
-------
--> 9 & 11
and -9 & -7

Answer by stanbon(75887)

product of two consecutive odd integers is 77 more than twice the larger,
Find the integers.
-------
Equation:
x(x+1) = 2(x+1)+77
-----
x^2+x = 2x+79
x^2-x-79 = 0
------
This quadratic does not have an integer solution.
==============
Cheers,
Stan H.

Answer by laurien56(2)

Geez Stan I came up with 9 and 11 because I did x(x+2) because they are two consecutive odd integers. Is that right?